Assume a Company Xs stock price today and cost of equity is consistent with the dividend discount model for constant growth HINT: Cost of Equity Dividend Yield + Capital Gain Yield. Use the information in the table below to answer problems 16-17. Any missing information will need to be solved for TIME (Years) 0 Shares Outstanding Book Value Prike Per Share Return On Equity Cost of Equity Long Term Growth Dividends Per Share 135.00 $6,750.00 80.00 12.0% 16. Company Xs dividend payout ratio is closest to: A. 0.3 B. 0.4 C. 0.5 D. 0.6 E. 0.7 17. The present value of Company Xs first dividend payment one year from now (T1) is closest to: A. $3.3.3 B. $3.40 C. $3.73 D. $3.81 E $3.83
Answer #1

1.
RoE=NI/Book Value=1350/6750=20%
Dividend payout ratio=1-growth rate/RoE=1-12%/20%=0.4

2.
cost of equity=(1350*0.4*(1+12%)/135)/80+12%=17.6%
Hence, present value of dividend=1350/135*0.4*(1+12%)/(1+17.6%)=3.81